# Performance Analysis of Physical Layer Security: From Far-Field to Near-Field

### v1

@@ IV-B1 UPW Model
Please refer to Appendix -G for more details. ∎
@@ IV-B3 NUSW Model
Under the NUSW model, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}\ll 1$ and $\lim_{M_{x},M_{z}\rightarrow\infty}G_{\mathsf{b}}^{\mathsf{N}}=\frac{A}{2d^{2}}$. Following the steps outlined in Appendix -G, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{P}=\lim_{M_{x},M_{z}\rightarrow\infty}\frac{2^{\mathsf{R}_{0}}-1}{\sigma^{-2}G_{\mathsf{b}}}=\frac{2\left(2^{\mathsf{R}_{0}}-1\right)d^{2}}{\sigma^{-2}A}$. ∎
@@
By comparing the results of the different channel models, we arrive at the following conclusion.
Remark 18.
Unlike FFC, NFC can achieve any desired secrecy rate for Bob, regardless of his location, provided there is sufficient transmit power. Moreover, the NUSW model is more realistic than the other models when investigating power scaling, as it adheres to the principle of energy conservation.
@@ V Numerical Results
In this section, numerical results are presented to demonstrate the secrecy performance in both the near field and far field, and to verify the accuracy of the analytical results. Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.125$ m, $d=\frac{\lambda}{2}$, $A=\frac{\lambda^{2}}{4\pi}$, $M_{x}=M_{z}=51$, $\overline{\gamma}=40$ dB, $\sigma^{2}=-10$ dB, $\mathsf{R}_{0}=1$ bps/Hz, $(\theta_{\mathsf{b}},\phi_{\mathsf{b}})=(\theta_{\mathsf{e}},\phi_{\mathsf{e}})=(\frac{\pi}{3},\frac{2\pi}{3})$, $r_{\mathsf{b}}=r$, and $r_{\mathsf{e}}=2r$ with $r=10$ m.
@@ V-A Secrecy Capacity
Figure 4 plots the secrecy capacity as a function of the number of transmit antennas for various values of $r$. It can be observed that for small values of $M$, the secrecy capacity under all channel models increases linearly with $\log M$. This occurs because, when the antenna number is small, the users can be treated as in the far-field region, where all models are sufficiently accurate. However, as $M$ grows larger, the variations in channel powers and phases across the array become pronounced. In such cases, the secrecy capacity under the UPW and USW models are overestimated due to neglect of these variations in wave propagation, and they will increase unboundedly with $M$, breaking the energy-conservation laws. This is in line with the statements in Remarks 4 and 6. In contrast, as $M$ increases, the secrecy capacity under the NUSW model are capped at finite values, aligning with the results in Remark 9 and justifying the superior accuracy and robustness of the NUSW model. Additionally, by examining the results for different values of $r$, we observe that the overestimation gap in secrecy capacity between the UPW/USW and NUSW models for large $M$ diminishes as the distance between the users and the BS increases. This occurs because, with greater distance, the users move toward the far-field region, where the near-field effects are mitigated.
@@
Figure 5 illustrates the secrecy channel capacity with respect to the distance of Eve for various channel models, where Bob and Eve are located in the same direction with $(\theta_{\mathsf{b}},\phi_{\mathsf{b}})=(\theta_{\mathsf{e}},\phi_{\mathsf{e}})=(\frac{\pi}{3},\frac{2\pi}{3})$. As can be seen, when Eve is closer to the BS than Bob, its channel condition poorer will be worse than Bob’s, and the secrecy capacity remains zero, making secure transmission unachievable. By contrast, secure transmission under NFC, i.e., under the USW and NUSW models, is always feasible except when Bob and Eve are at the exact same location. These behaviors are consistent with the insights provided in Remarks 3 and 8. Moreover, NFC consistently achieves higher secrecy capacity than FFC, even when Bob is farther from the BS than Eve. This advantage is attributed to the enhanced distance resolution offered by the spherical-wave model, which enables more precise focusing of signal energy on Bob while minimizing leakage to Eve. These observations underscore the superior secrecy performance of NFC.
@@
Figure 9: Minimum required power versus $r_{\mathsf{e}}$ when $r_{\mathsf{b}}=10$ m.
V-B Minimum Required Power
In Figure 8, we present the minimum required power as a function of the antenna number $M$. Initially, for smaller values of $M$, the minimum required power under all models approximately decreases linearly with $\log M$. However, as $M$ increases, expanding the near-field region, the UPW and USW models become less accurate. Specifically, the minimum required power under the UPW/USW models continues to decrease and approaches zero, which is unrealistic in practical scenarios. In contrast, for the more accurate NUSW model, the minimum required power converges to a lower bound greater than zero, reflecting a more realistic scenario. These observations corroborate the insights provided in Remarks 15, 16 and 17.
Figure 9 shows the minimum required power as a function of Eve’s distance from the BS for different target secrecy rates. It is worth noting that, when Bob and Eve are in the same direction, under FFC, the target secrecy rate is only achievable when $r_{\mathsf{e}}>2^{\mathsf{R}_{0}/2}r_{\mathsf{b}}$. In contrast, NFC can achieve the target rate whenever $r_{\mathsf{e}}\neq r_{\mathsf{b}}$, in line with Remark 18. Furthermore, for a given $\mathsf{R}_{0}$, the required transmit power for FFC is always higher than that of NFC, demonstrating the advantages of NFC in enhancing physical layer security.
@@ VI Conclusion
$\mathsf{C}\approx\log_{2}\left(\frac{\overline{\gamma}_{\mathsf{b}}G_{\mathsf{b}}\overline{\gamma}_{\mathsf{e}}G_{\mathsf{e}}\left(1-\rho\right)}{\overline{\gamma}_{\mathsf{e}}G_{\mathsf{e}}}\right)=\log_{2}\left(\overline{\gamma}_{\mathsf{b}}G_{\mathsf{b}}\left(1-\rho\right)\right),$
@@
To calculate $\delta_{i}$, we next consider different cases.
@@
$\delta_{i}=\left|\sum\nolimits_{m_{i}=-\tilde{M}_{i}}^{\tilde{M}_{i}}{\mathrm{e}^{\mathrm{j}b_{i}m_{i}}}\right|^{2}=\left|\frac{\sin\left(M_{i}b_{i}/2\right)}{\sin\left(b_{i}/2\right)}\right|^{2}.$
@@
$\overset{\left(a\right)}{\approx}\!M_{i}^{2}\left|\int_{-\frac{1}{2}}^{\frac{1}{2}}\!{\mathrm{e}^{\mathrm{j}\vartheta_{i}x^{2}}\mathrm{d}}x\right|^{2}\!\overset{\left(b\right)}{=}\frac{\pi}{\left|a_{i}\right|}\left|\mathrm{erf}\!\bigg{(}\frac{\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}M_{i}\sqrt{\left|a_{i}\right|}}{2}\bigg{)}\right|^{2},$
where approximation $(a)$ follows from the concept of definite integral for large $T$, equality $(b)$ follows from the integral identity [22, Eq. (2.33.3)].
@@
=$\overset{\left(a\right)}{=}\!\frac{\pi}{4\left|a_{i}\right|}\left|\mathrm{erf}\bigg{(}\frac{a_{i}M_{i}-b_{i}}{2\sqrt{\left|a_{i}\right|}}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\bigg{)}+\mathrm{erf}\bigg{(}\frac{a_{i}M_{i}+b_{i}}{2\sqrt{\left|a_{i}\right|}}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\bigg{)}\right|^{2},$
where $(a)$ follows from the integral identity [22, Eq. (2.33.3)].
@@
Based on (6b) and (10), the channel gains under the NUSW model can be written as follows:
@@
$f\left(x,z\right)\triangleq(x^{2}+z^{2}-2\Phi_{k}x-2\Omega_{k}z+1)^{-\frac{3}{2}}$ (A11)
in the area $\mathcal{A}=\left\{\left(x,z\right)\mid-\frac{M_{i}\varepsilon_{k}}{2}\leq i\leq\frac{M_{i}\varepsilon_{k}}{2},i\in\left\{x,z\right\}\right\}$ which is then partitioned into $M_{x}M_{z}$ sub-rectangles, each with equal area $\varepsilon_{k}^{2}$. Since $\varepsilon_{k}\ll 1$, we have $f\left(x,z\right)\approx f\left(m_{x}\varepsilon_{k},m_{z}\varepsilon_{k}\right)$ for $\forall\left(x,z\right)\in\left\{\left(x,z\right)\mid\left(m_{i}-\frac{1}{2}\right)\varepsilon_{k}\leq i\leq\left(m_{i}+\frac{1}{2}\right)\varepsilon_{k},i\in\left\{x,z\right\}\right\}$.
@@
$\sum\nolimits_{m_{x},m_{z}}{f\left(m_{x}\varepsilon_{k},m_{z}\varepsilon_{k}\right)\varepsilon_{k}^{2}}\approx\iint_{\mathcal{A}}{f\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$ (A12)
@@
$\begin{split}G_{k}&=\frac{A^{2}\Psi_{k}}{4\pi d^{2}}\int_{-\frac{M_{z}\varepsilon_{k}}{2}}^{\frac{M_{z}\varepsilon_{k}}{2}}{\int_{-\frac{M_{x}\varepsilon_{k}}{2}}^{\frac{M_{x}\varepsilon_{k}}{2}}}f(x,z)\mathrm{d}x\mathrm{d}z.\end{split}$ (A13)
We can calculate the inner integral with the aid of [22, Eq. (2.264.5)] and the outer integral with the aid of [22, Eq. (2.284.5)], which yields the results of (39).
@@
-E Proof of Lemma 2
@@
$\rho\triangleq\left|{\mathrm{erf}\left(\sqrt{\pi}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\Upsilon\right)}/{(2\Upsilon)}\right|^{4},$ (A15)
where $\Upsilon=\sqrt{\frac{Md^{2}}{4\lambda}|{1}/{r_{\mathsf{b}}}-{1}/{r_{\mathsf{e}}}|}$. It can be clearly shown from the numerical results that $\left|{\mathrm{erf}\left(\sqrt{\pi}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\Upsilon\right)}/{(2\Upsilon)}\right|^{4}\geq\frac{1}{2}$ if and only if $\Upsilon\leq 0.79$. Consequently, by defining $\Upsilon_{3\mathrm{dB}}\triangleq 0.79$, we have $1-\rho\leq\frac{1}{2}$ when
$\sqrt{{Md^{2}}/({4\lambda})\left|{r_{\mathsf{b}}^{-1}}-{r_{\mathsf{e}}^{-1}}\right|}\leq\Upsilon_{3\mathrm{dB}}.$ (A16)
@@
-F Proof of Theorem 5
@@
$\mathsf{P}={(2^{\mathsf{R}_{0}}-1)}/({\sigma^{-2}G_{\mathsf{b}}^{\mathsf{P}}-2^{\mathsf{R}_{0}}\sigma^{-2}G_{\mathsf{e}}^{\mathsf{P}}}).$ (A27)

### v2

@@ IV-B1 UPW Model
Please refer to Appendix -H for more details. ∎
@@ IV-B3 NUSW Model
Under the NUSW model, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}\ll 1$ and $\lim_{M_{x},M_{z}\rightarrow\infty}G_{\mathsf{b}}^{\mathsf{N}}=\frac{A}{2d^{2}}$. Following the steps outlined in Appendix -H, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{P}=\lim_{M_{x},M_{z}\rightarrow\infty}\frac{2^{\mathsf{R}_{0}}-1}{\sigma^{-2}G_{\mathsf{b}}}=\frac{2\left(2^{\mathsf{R}_{0}}-1\right)d^{2}}{\sigma^{-2}A}$. ∎
@@
By comparing the results of the different channel models, we can find that the NUSW model is more realistic than the other models when investigating power scaling, as it adheres to the principle of energy conservation.
@@ V Numerical Results
In this section, numerical results are presented to demonstrate the secrecy performance in both the near field and far field, and to verify the accuracy of the analytical results. Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.125$ m, $d=\frac{\lambda}{2}$, $A=\frac{\lambda^{2}}{4\pi}$, $M_{x}=M_{z}=51$, $\overline{\gamma}=40$ dB, $\sigma^{2}=-10$ dB, $\mathsf{R}_{0}=1$ bps/Hz, $(\theta_{\mathsf{b}},\phi_{\mathsf{b}})=(\theta_{\mathsf{e}},\phi_{\mathsf{e}})=(\frac{\pi}{3},\frac{2\pi}{3})$, $r_{\mathsf{b}}=r$ and $r_{\mathsf{e}}=2r$ with $r=10$ m, and $T=100$.
@@ V-A Secrecy Capacity
Figure 5 plots the secrecy capacity as a function of the number of transmit antennas for various values of $r$. It can be observed that for small values of $M$, the secrecy capacities under all channel models increase linearly with $\log M$. This occurs because, when the antenna number is small, the users can be treated as in the far-field region, where all models are sufficiently accurate. However, as $M$ grows larger, the variations in channel powers and phases across the array become pronounced. In such cases, the secrecy capacities under the UPW and USW models are overestimated due to neglect of these variations in wave propagation, and they will increase unboundedly with $M$, breaking the energy-conservation laws. This is in line with the statements in Remarks 4 and 6. In contrast, as $M$ increases, the secrecy capacity under the NUSW model is capped at finite values, aligning with the results in Remark 9 and justifying the superior accuracy and robustness of the NUSW model.
@@
(a) NUSW.
(b) USW.
(c) UPW.
Figure 7: Normalized secrecy capacity versus small angular difference between Bob and Eve.
Figure 6 illustrates the secrecy channel capacity with respect to the distance of Eve for various channel models, where Bob and Eve are located in the same direction with $(\theta_{\mathsf{b}},\phi_{\mathsf{b}})=(\theta_{\mathsf{e}},\phi_{\mathsf{e}})=(\frac{\pi}{3},\frac{2\pi}{3})$. As can be seen, when Eve is closer to the BS than Bob, its channel condition poorer will be worse than Bob’s, and the far-field secrecy capacity remains zero, making secure transmission within the far field unachievable. By contrast, secure transmission under NFC is always feasible except when Bob and Eve are at the exact same location. These behaviors are consistent with the insights provided in Remarks 3 and 8, underscoring the superior secrecy performance of NFC. Moreover, we can observe that NFC consistently achieves higher secrecy capacity than FFC, even when Bob is farther from the BS than Eve. This advantage is attributed to the enhanced distance resolution offered by the spherical-wave model, which enables more precise focusing of signal energy on Bob while minimizing leakage to Eve. However, this gap decreases as $r_{\mathsf{e}}$ increase. This is because the near-field effect is more pronounced at short distances and diminishes as Bob or Eve moves toward the far-field region.
While we find that the secrecy capacity for co-directional Bob and Eve improves under NFC compared to the FF case, it is nearly impossible for them to be positioned in exactly the same direction in real-word scenarios, typically only similarly. Therefore, it would be important to explore how small angular perturbations between Bob and Eve impact the secrecy capacity. By letting $(\theta_{\mathsf{e}},\phi_{\mathsf{e}})=(\theta_{\mathsf{b}}+\Delta\theta,\phi_{\mathsf{b}}+\Delta\phi)$ with $(\Delta\theta,\Delta\phi)$ denoting the small angular perturbations, Fig. 7 illustrates how secrecy capacity fluctuates with changes in $(\Delta\theta,\Delta\phi)$. The results show that the capacities for the NUSW and USW models remain high regardless of the angle differences. On the other hand, while the secrecy capacity under the UPW model is comparable to NFC when Bob and Eve are in different directions, it significantly deteriorates when their angular difference is minimal. These observations confirm the superior performance of NFC in scenarios where Eve is aligned with Bob or even with small angular deviations.
@@
Figure 10: Minimum required power versus $\mathsf{R}_{0}$.
V-B Minimum Required Power
Fig. 10 illustrates the minimum required power for various channel models to achieve different target secrecy rates when Bob and Eve are located in the same direction. It can be observed that the UPW model requires more transmission power than the USW and NUSW models to achieve a same target secrecy rate, with the disparity increasing as $\mathsf{R}_{0}$ grows. Notably, when the target secrecy rate exceeds the threshold $\mathsf{R}_{0}>2\log_{2}\frac{r_{\mathsf{e}}}{r_{\mathsf{b}}}$, the UPW model becomes incapable of achieving it, regardless of the available transmission power. Conversely, the near-field model can attain arbitrarily high secrecy rates, given sufficient transmission power. These findings are consistent with the discussions in Remark 15. Additionally, Figure 11 shows the minimum required power as a function of Eve’s distance from the BS. It is worth noting that the under FFC, a target secrecy rate is only achievable when $r_{\mathsf{e}}>2^{\mathsf{R}_{0}/2}r_{\mathsf{b}}$, whereas NFC can achieve the target rate as long as $r_{\mathsf{e}}\neq r_{\mathsf{b}}$. The above observations demonstrate the advantages of NFC in enhancing physical layer security.
Figure 11: Minimum required power versus $r_{\mathsf{e}}$ with $r_{\mathsf{b}}=10$ m.
@@ V-B Minimum Required Power
In Figure 12, we present the minimum required power as a function of the antenna number $M$. Initially, for smaller values of $M$, the minimum required power under all models approximately decreases linearly with $\log M$. However, as $M$ increases, expanding the near-field region, the UPW and USW models become less accurate. Specifically, the minimum required power under the UPW/USW models continues to decrease and approaches zero, which is unrealistic in practical scenarios. In contrast, for the more accurate NUSW model, the minimum required power converges to a lower bound greater than zero, reflecting a more realistic scenario. These observations corroborate the insights provided in Remarks 16, 17 and 18.
@@ VI Conclusion
$\mathsf{C}\approx\log_{2}\!\left(\frac{\overline{\gamma}_{\mathsf{b}}G_{\mathsf{b}}\overline{\gamma}_{\mathsf{e}}G_{\mathsf{e}}\left(1-\rho\right)}{\overline{\gamma}_{\mathsf{e}}G_{\mathsf{e}}}\right)\!=\log_{2}\!\left(\overline{\gamma}_{\mathsf{b}}G_{\mathsf{b}}\left(1\!-\!\rho\right)\right),$ (A8)
@@
Since Bob and Eve are at different locations, $a_{i}$ and $b_{i}$ cannot be zero at the same time. To calculate $\delta_{i}$, we consider different cases as follows.
@@
$\delta_{i}=\left|\sum\nolimits_{m_{i}=-\tilde{M}_{i}}^{\tilde{M}_{i}}{\mathrm{e}^{\mathrm{j}b_{i}m_{i}}}\right|^{2}=\left|\frac{\sin\left(M_{i}b_{i}/2\right)}{\sin\left(b_{i}/2\right)}\right|^{2}.$ (A11)
@@
$\!\overset{\left(a\right)}{\approx}\!M_{i}^{2}\left|\int_{-\frac{1}{2}}^{\frac{1}{2}}\!{\mathrm{e}^{\mathrm{j}\vartheta_{i}x^{2}}\mathrm{d}}x\right|^{2}\!\!\overset{\left(b\right)}{=}\frac{\pi}{\left|a_{i}\right|}\left|\mathrm{erf}\bigg{(}\frac{\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}M_{i}\sqrt{\left|a_{i}\right|}}{2}\bigg{)}\right|^{2}\!\!,$ (A12)
where approximation $(a)$ follows from the concept of definite integral for large $T$, equality $(b)$ follows from the integral identity [32, Eq. (2.33.3)].
@@
=$\overset{\left(a\right)}{=}\!\frac{\pi}{4\left|a_{i}\right|}\left|\mathrm{erf}\bigg{(}\frac{a_{i}M_{i}-b_{i}}{2\sqrt{\left|a_{i}\right|}}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\bigg{)}+\mathrm{erf}\bigg{(}\frac{a_{i}M_{i}+b_{i}}{2\sqrt{\left|a_{i}\right|}}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\bigg{)}\right|^{2},$ (A13)
where $(a)$ follows from the integral identity [32, Eq. (2.33.3)].
@@
Based on (6) and (10), the channel gain for user $k\in\{\mathrm{b},\mathrm{e}\}$ under the NUSW model can be written as follows:
@@
$f_{k}\left(x,z\right)\triangleq(x^{2}+z^{2}-2\Phi_{k}x-2\Omega_{k}z+1)^{-\frac{3}{2}}$ (A15)
in the area $\mathcal{A}=\left\{\left(x,z\right)\mid-\frac{M_{i}\varepsilon_{k}}{2}\leq i\leq\frac{M_{i}\varepsilon_{k}}{2},i\in\left\{x,z\right\}\right\}$ which is then partitioned into $M_{x}M_{z}$ sub-rectangles, each with equal area $\varepsilon_{k}^{2}$. Since $\varepsilon_{k}\ll 1$, we have $f_{k}\left(x,z\right)\approx f_{k}\left(m_{x}\varepsilon_{k},m_{z}\varepsilon_{k}\right)$ for $\forall\left(x,z\right)\in\left\{\left(x,z\right)\mid\left(m_{i}-\frac{1}{2}\right)\varepsilon_{k}\leq i\leq\left(m_{i}+\frac{1}{2}\right)\varepsilon_{k},i\in\left\{x,z\right\}\right\}$.
@@
$\sum\nolimits_{m_{x},m_{z}}{f_{k}\left(m_{x}\varepsilon_{k},m_{z}\varepsilon_{k}\right)\varepsilon_{k}^{2}}\approx\iint_{\mathcal{A}}{f_{k}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$ (A16)
@@
$G_{k}=\frac{A\Psi_{k}}{4\pi d^{2}}\int_{-\frac{M_{z}\varepsilon_{k}}{2}}^{\frac{M_{z}\varepsilon_{k}}{2}}{\int_{-\frac{M_{x}\varepsilon_{k}}{2}}^{\frac{M_{x}\varepsilon_{k}}{2}}}f_{k}(x,z)\mathrm{d}x\mathrm{d}z.$ (A17)
We can calculate the inner integral with the aid of [32, Eq. (2.264.5)] and the outer integral with the aid of [32, Eq. (2.284.5)], which yields the results of (41).
@@
-E Proof of Corollary 13
For the ULA case, the channel gains under the NUSW model can be written as follows:
$G_{k}=\frac{A\Psi_{k}}{4\pi r_{k}^{2}}\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}(m_{z}^{2}\varepsilon_{k}^{2}-2m_{z}\varepsilon_{k}\Omega_{k}+1)^{-\frac{3}{2}}.$ (A19)
[…]
$G_{k}=\frac{A\Psi_{k}}{4\pi d^{2}}\int_{-\frac{M\varepsilon_{k}}{2}}^{\frac{M\varepsilon_{k}}{2}}f^{\prime}_{k}(z)\mathrm{d}z,$ (A20)
The integral can be calculated with the aid of [32, Eq. (2.264.5)], leading to the results of (47).
Following the steps to obtain (A20), the channel correlation factor can be expressed as follows:
$\rho=\frac{A^{2}\Psi_{\mathsf{b}}\Psi_{\mathsf{e}}}{16\pi^{2}G_{\mathsf{b}}^{\mathsf{N}}G_{\mathsf{e}}^{\mathsf{N}}r_{\mathsf{b}}^{2}r_{\mathsf{e}}^{2}\varepsilon_{\mathsf{b}}^{4}}\left|\int_{-\frac{M\varepsilon_{\mathsf{b}}}{2}}^{\frac{M\varepsilon_{\mathsf{b}}}{2}}{{g^{\prime}_{1}\left(z\right)g^{\prime}_{2}\left(z\right)}\mathrm{d}z}\right|^{2}.$ (A21)
By applying the Chebyshev-Gauss quadrature rule to the above integral, we can obtain the results of (48).
-F Proof of Lemma 2
@@
$\rho\triangleq\left|\frac{\mathrm{erf}\left(\sqrt{\pi}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\Upsilon\right)}{2\Upsilon}\right|^{4},$ (A22)
where $\Upsilon=\sqrt{\frac{Md^{2}}{4\lambda}|{r_{\mathsf{b}}^{-1}}-{r_{\mathsf{e}}^{-1}}|}$. It can be clearly shown from the numerical results given in Fig. 13 that $\left|\frac{\mathrm{erf}\left(\sqrt{\pi}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\Upsilon\right)}{2\Upsilon}\right|^{4}\geq\frac{1}{2}$ if and only if $\Upsilon\leq 0.79$. Consequently, by defining $\Upsilon_{3\mathrm{dB}}\triangleq 0.79$, we have $1-\rho\leq\frac{1}{2}$ when
$\sqrt{\frac{Md^{2}}{4\lambda}\left|{r_{\mathsf{b}}^{-1}}-{r_{\mathsf{e}}^{-1}}\right|}\leq\Upsilon_{3\mathrm{dB}}.$ (A23)
@@
Figure 13: Numerical results of $\left|{\mathrm{erf}\left(\sqrt{\pi}\mathrm{e}^{\mathrm{j}\frac{\pi}{4}}\Upsilon\right)}/{(2\Upsilon)}\right|^{4}$.
-G Proof of Theorem 5
@@
$\mathsf{P}=\frac{2^{\mathsf{R}_{0}}-1}{\sigma^{-2}G_{\mathsf{b}}^{\mathsf{P}}-2^{\mathsf{R}_{0}}\sigma^{-2}G_{\mathsf{e}}^{\mathsf{P}}}.$ (A34)